Elimination:How do you figure out the problems:


3x + 2y = 9 -x + 3y = 8

5X +3Y = 1.5
-8x - 2y = 20

Answer by checkley71(8403)

3X+2Y=9 NOW MULTIPLY THE OTHER EQUATION BY 3 & ADD THEM
-3X+9Y=24
--------------------------------
11Y=33
Y=33/11
Y=3 ANSWER.
3X+2*3=9
3X+6=9
3X=9-6
3X=3
X=3/3
X=1 ANSWER.